OpenStudy (anonymous):

\int\limits_{}^{}x ^{2}(x ^{3}+5)^{9} dx

OpenStudy (anonymous):

u=x^3+5 du=3x^2 \[\int\limits 1/3u ^{9} du\]=1/30*(x^3+5)^10